# Constrained LTL Specification Learning from Examples

### v1

@@ V-D Learning Objective Encoding
The following code allows $\textnormal{Alloy}^{\textnormal{Max}}$ to generate an LTL formula satisfying the sample.
@@
3} for /* max number of DAG node */ DAGNode
Line 2 defines that for any positive trace $t\in P$, $(t,root,0)\in val$, i.e., the learned formula is true on trace $t$ from time 0. In contrast, for any negative trace $t\in N$, $(t,root,0)$ is not in $val$ (line 3). Finally, on line 4, the user provides the max number of DAG nodes allowed for a particular problem.
@@ V-F Optimization Objective
Similar to the syntactic constraints, $\textnormal{Alloy}^{\textnormal{Max}}$ supports all the optimization operators for defining optimization objective $\Psi$. Specifically, it has the following mappings:
@@ VI-A Solving $\textnormal{Alloy}^{\textnormal{Max}}$ with MaxSAT
We briefly explain how Alloy and $\textnormal{Alloy}^{\textnormal{Max}}$ solve a problem using a SAT and a MaxSAT solver; however, interested readers should refer to the original papers for more details [15, 29, 30]. A relation in Alloy is translated into a matrix of Boolean variables – each of which is true if and only if the tuple represented by this particular variable is in the relation; and a relational expression is represented by operations over one or more Boolean matrices.
@@ VI-B Correctness
Our approach is sound but complete only up to the bound of the user-provided max number of DAG nodes. A correctness proof can be found in the Appendix.
@@ VI-C Quality of Solutions
Disable reusing in DAG: The DAG encoding was designed for improving performance. When minimizing against a DAG, it guarantees a minimal number of distinct sub-formulas. However, this does not always produce a minimal LTL formula in its syntax tree size where repeating sub-formulas are also counted. For example, consider $\mathbf{F}p\lor\mathbf{F}\mathbf{G}p$ and $\mathbf{F}\mathbf{G}p\lor\mathbf{F}\mathbf{G}p$. The latter one has a bigger syntax tree size but a smaller DAG size as the sub-formula $\mathbf{F}\mathbf{G}p$ is reused [8]. Such behavior might reduce the usefulness of the tool for certain problems and can be disabled with the following constraint:
@@
Negation normal form (NNF): The use of negation may result in some equivalent but less-readable formulas, e.g., $\neg\mathbf{G}(x)$ versus $\mathbf{F}(\neg x)$ where the latter may be more readable and preferable. NNF is also helpful in automated theorem proving [31]. Thus, we can enforce that negation should only be applied to atomic propositions.
@@
This is an inexhaustive list of problem-independent constraints that we find helpful. Other examples include requiring $\phi$ being in conjunction normal form (CNF) or disjunction normal form (DNF).
@@ VII Use Cases
In this section, we present three use cases to demonstrate the need for syntactic constraints and/or optimization objectives when performing LTL learning in real-world problems. These use cases also show how tool builders can leverage ATLAS as a back-end LTL learner to perform various specification-based tasks, such as repair and invariant weakening.
@@ VII-A Specification Mining
This use case demonstrates the usefulness of custom constraints and enumeration. This case study is motivated by a case of voter fraud in Clay County, Kentucky, in which corrupt officials manipulated a flaw in the system to “flip” votes [32]. Figure 7 shows an abstract model of it as described in [33]. A voter enters the booth, enters her password, selects a candidate, votes, and then finally confirms before leaving the booth.
We generate example traces based on this model, manually check their correctness, and mark them as positive or negative. Specifically, a negative trace indicating a vulnerability of this system is: a voter leaves the booth before confirming her vote; then, a corrupt official enters the booth and uses the back button to select another candidate, and confirms the choice.
@@
We want to learn a safety invariant to ensure election integrity. Particularly, the invariant should be in $\mathbf{G}\phi$ where $\phi$ should not contain any temporal operators. These constraints can be expressed as:
@@
In addition, our technique supports enumerating solutions that satisfy the given constraints, in the order of the formula size (or other custom optimization objectives). This is often helpful in finding variants of formulas. Specifically, in this case study, the enumeration helps us find two safety invariants:
@@ VII-B Specification Repair from Demonstrations
In this use case, we demonstrate how custom optimization objectives can be helpful in learning. The use case is motivated by the field of autonomous agent interpretability using LTL specifications[34]. Consider, for example, the setup in Figure 8a where a robotic arm’s end effector has to be actuated to achieve some tasks. The designer provides an initial LTL specification: $\mathbf{F}(green)\land\mathbf{G}(\neg red)$; i.e., reaching the green region and avoiding the red region.
However, after a controller has been generated, the designer decides to add additional demonstrations to help “repair” the LTL specification to better reflect the needs. For example, consider demonstrations satisfying the above LTL formula but also visiting the blue region. Thus, the current specification needs to be modified to factor in this change. The developer can mark the expected demonstration traces as positive (i.e., the blue region reached) and mark the others as negative. Moreover, the modification should be minimized. We pose this as an LTL learning problem while minimizing the editing distance, i.e., minimizing the removal of existing sub-formulas as well as the addition of new sub-formulas. The encoding of this minimal modification problem is defined as follows:
@@
Lines (8)-(10) define the original formula. Line (11) minimizes the editing distance by using the $max$ operator (with priority 2). The learning process will first retain as many sub-formulas as possible from the old specification and then minimize the total size of the final formula. With these constraints, our approach can learn the ideal specification: $\mathbf{F}(green)\land\mathbf{G}(\neg red)\land\mathbf{F}(blue)$, where $blue$ stands for reaching the blue region.
@@ VII-C Invariant Weakening
For instance, Figure 8b is a radiation therapy machine similar to Therac-25 [40] as described in [41]. The machine has two modes: Electron Beam mode and X-ray mode. A spreader must be inserted during the X-ray mode to attenuate the effect of the high-power X-ray beam and limit possible overdose. Therefore, we can define a safety invariant $P$: $\mathbf{G}(\textsf{XrayMode}\rightarrow\textsf{SpreaderIn})$.
@@
In general, for a safety invariant in $\mathbf{G}(\phi\rightarrow\psi)$ where $\phi$ and $\psi$ are Boolean formulas. To achieve weakening, one can learn an invariant $\mathbf{G}(\phi\land\phi^{\prime}\rightarrow\psi\lor\psi^{\prime})$ where $\phi\land\phi^{\prime}$ is a Boolean formula in CNF and $\psi\lor\psi^{\prime}$ is a Boolean formula in DNF. In other words, an invariant can be weakened by (1) adding more conditions (as conjunctions) in the assumption, or (2) adding new acceptable conditions (as disjunctions) in the guarantee. This weakening objective can be expressed through the following constraints:
@@
Specifically, line (12) defines the structure $\mathbf{G}(\phi\rightarrow\psi)$. Line (13) defines $\phi$ and $\psi$ contain only ,,$\land$,$\lor$,$\neg$ , ∨ , ¬, and $AP$. Lines (14)-(16) define $\phi$ is in CNF and $\psi$ is in DNF. Lines (17)-(18) define $\phi$ may add conjunctions to XrayMode, and $\psi$ may add disjunctions to SpreaderIn. With these constraints, we can learn the ideal weakened formula as described above.
@@ VIII Evaluation
    RQ1: How does our approach perform in comparison to the state-of-the-art tool, Flie [9], over unconstrained LTL learning problems?
@@
    RQ2: For problems with structural constraints and/or optimization objectives as described in our case studies, how does our approach perform in comparison to the state-of-the-art tools?
For RQ1, the metric we are concerned with is the total time taken to find the first satisfying formula. For RQ2, we evaluate performance through two key metrics: (1) the number of constrained problems solved and (2) the time taken to solve them. In particular, even though existing tools cannot encode constraints such as ours, they can be configured to enumerate all solutions until a solution satisfying the constraints is found; the goal here is to compare how long it takes for ATLAS and the baselines in finding an ideal solution.
@@ VIII-A Experimental Setup
For fair comparison, we compare ATLAS against the state-of-the-art tools satisfying the following two criteria: (1) no restrictions on the type of formulas that can be learned and (2) the guarantee of finding minimal formulas.
@@ VIII-A2 RQ2
For constrained problems, we first compare our tool against Flie, which is configured to enumerate solutions until an expected formula is found. We also compare against LTLSketcher [8] that can learn a formula given a user-defined template. It considers three types of placeholders in a template: (1) any LTL sub-formula, (2) any unary operator, and (3) any binary operator. However, it cannot express all our expected constraints and does not support enumeration either. Thus, we test whether it finds an expected formula in one run with the closest template regarding our required constraints. If it does not, we count it as a timeout. Also note that, LTLSketcher is not used in RQ1 as it builds on Flie, and without user-provided templates, it reduces to Flie.
@@ VIII-B Results
Moreover, Table I shows the number of solved problems and the average solving time for Flie, LTLSketcher, and ATLAS. Our tool solves the most problems and also has the minimal average solving time. On average, we solve about 3.2x as many problems as Flie and 8.7x as many as LTLSketcher. Also, our tool is about 3x faster than both Flie and LTLSketcher.
Although ATLAS also failed to solve some problems due to timeout, it is better than Flie and LTLSketcher in that: Flie needs to enumerate a large, unconstrained search space which leads to timeout more often. LTLSketcher failed due to the lack of expressiveness (in defining the expected constraints) and enumeration capability. Thus, we show that strong expressive power and enumeration capability are both critical to LTL learning, and our tool supports both. We can solve all problems that Flie and LTLSketcher can without the loss of generality and much performance overhead.
@@ IX Related Work
Before Neider’s approach, techniques for mining LTL, such as [13, 4, 12], rely on patterns or templates. For example, given a template $\mathbf{G}(x\rightarrow\mathbf{F}y)$, Texada [4] infers specifications by substituting $x$ and $y$ with atomic events from the given traces (and does not learn from negative traces). Compared to our approach, Texada lies on a different trade-off point between expressiveness and efficiency: Although ATLAS gives the user more control over the properties of LTL expressions to be learned, Texada is able to mine from much larger traces (e.g., thousands of events). More recently, Lutz et al. present LTLSketcher [8], where they consider three types of substitutions (i.e., any formula, unary operator, and binary operator) and propose a SAT-based method to find valid substitutions efficiently. Although they have improved expressiveness compared to Texada, they cannot express many patterns (e.g., CNF and DNF) as demonstrated in our use cases. In addition, none of these tools support custom optimization objectives.
Researchers have also explored different variations of the learning problem [47]. Gaglione et al. [48] presents an approach using MaxSAT to learn from noisy data in the sense that not all the examples must be satisfied or falsified. Although $\textnormal{Alloy}^{\textnormal{Max}}$ also uses MaxSAT, we cannot solve this type of problems because $\textnormal{Alloy}^{\textnormal{Max}}$ cannot directly control the weights of clauses in a MaxSAT problem. Roy et al. [49] present an approach for learning from positive-only examples. While our tool also supports such a problem, we can only constrain and optimize against the syntactic structure of the solutions whereas they can guarantee semantic minimality (i.e., a minimal set of behavior).
@@ X Conclusion and Future Work
We have proposed the constrained LTL learning problem as a generalization of LTL learning from examples, and demonstrated the applicability and performance of ATLAS, our prototype implementation tool. The scalability of our tool is limited by the underlying MaxSAT solver, and an interesting future work is to investigate an approach that uses a different learning method (e.g., based on machine learning). We also plan to investigate methods or heuristics for further reducing the space of possible solutions (e.g., an encoding that eliminates redundant, equivalent LTL expressions). Moreover, we plan to investigate additional use cases for ATLAS (such as invariant synthesis for distributed systems).
Our approach improves the expressive power of LTL learning tools. However, crafting constraints in FOL can be challenging for non-experts in formal logic. Thus, we also plan to improve the usability of LTL learning by building on ATLAS (e.g., through a DSL front-end or an LLM that converts natural language to FOL). In addition, ATLAS and other SOTA tools require users to provide examples. Future work could explore integrating external methods (e.g., model checkers or test generation tools) to automate the generation of examples.
@@ Acknowledgment
This work was supported in part by the NSF awards 1918140, 2144860 and 2319317, DoD contract FA8702-15-D-0002, DARPA contract FA87501620042, and the CAMELOT project (POCI-01-0247-FEDER045915) from the Portuguese Science and Technology Foundation. Any views, opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the organizations.

### v2

@@ V-D Learning Objective Encoding
The following code allows $\textnormal{Alloy}^{\textnormal{Max}}$ to generate an LTL formula that satisfies the sample.
@@
3} for /* max number of DAG nodes */ DAGNode
Line 2 defines that for any positive trace $t\in P$, $(t,root,0)\in val$, i.e., the learned formula is true on trace $t$ from time 0. In contrast, for any negative trace $t\in N$, $(t,root,0)$ is not in $val$ (line 3). Finally, on line 4, the user provides the maximum number of DAG nodes allowed for a particular problem.
@@ V-F Optimization Objective
Similar to the syntactic constraints, $\textnormal{Alloy}^{\textnormal{Max}}$ supports all the optimization operators for defining objective $\Psi$. Specifically, it has the following mappings:
@@ VI-A Solving $\textnormal{Alloy}^{\textnormal{Max}}$ with MaxSAT
We briefly explain how Alloy and $\textnormal{Alloy}^{\textnormal{Max}}$ solve a problem using a SAT and a MaxSAT solver, respectively; interested readers should refer to the original papers for more details [15, 29, 30]. A relation in Alloy is translated into a matrix of Boolean variables – each of which is true if and only if the tuple represented by this particular variable is in the relation; and a relational expression is represented by operations over one or more Boolean matrices.
@@ VI-B Correctness
Our approach is sound but complete only up to the user-provided bound on the maximum number of DAG nodes. A correctness proof can be found in the Appendix.
@@ VI-C Quality of Solutions
Disable reusing in DAG: The DAG encoding was designed to improve performance. When minimizing against a DAG, it guarantees a minimal number of distinct sub-formulas. However, this does not always produce a minimal LTL formula in its syntax tree size where repeating sub-formulas are also counted. For example, consider $\mathbf{F}p\lor\mathbf{F}\mathbf{G}p$ and $\mathbf{F}\mathbf{G}p\lor\mathbf{F}\mathbf{G}p$. The latter one has a bigger syntax tree size but a smaller DAG size as the sub-formula $\mathbf{F}\mathbf{G}p$ is reused [8]. Such behavior might reduce the usefulness of the tool for certain problems and can be disabled with the following constraint:
@@
Negation normal form (NNF): The use of negation may result in some equivalent but less readable formulas, e.g., $\neg\mathbf{G}(x)$ versus $\mathbf{F}(\neg x)$ where the latter may be more readable and preferable. NNF is also helpful in automated theorem proving [31]. Thus, we can enforce that negation should only be applied to atomic propositions.
@@
This is an inexhaustive list of problem-independent constraints that we find helpful. Other examples include requiring $\phi$ to be in conjunction normal form (CNF) or disjunction normal form (DNF).
@@ VII Use Cases
In this section, we present three use cases to demonstrate the need for syntactic constraints and/or optimization objectives. These use cases also show how tool builders can leverage ATLAS as a back-end LTL learner to perform various specification-based tasks, such as specification repair and invariant weakening.
@@ VII-A Specification Mining
This use case demonstrates the usefulness of custom constraints and enumeration. This case study is inspired by a voter fraud incident in Kentucky, USA, where corrupt officials manipulated a flaw in the system to “flip” votes [32]. Figure 7 shows an abstract model of the system as described in [33]: A voter enters the booth, enters a password, selects a candidate, votes, and finally confirms the vote before leaving the booth.
We generate example traces based on this model, manually check their correctness, and mark them as positive or negative. Specifically, a negative trace indicating a vulnerability of this system is: A voter leaves the booth before confirming their vote; then, a corrupt official enters the booth and uses the back button to select another candidate and confirms the choice.
@@
We want to learn a safety invariant to ensure election integrity. Particularly, the invariant should be of the form $\mathbf{G}\phi$, where $\phi$ should not contain any temporal operators. These constraints can be expressed as:
@@
In addition, our technique supports enumerating solutions that satisfy the given constraints, in the order of the formula size (or other custom optimization objectives). This is often helpful in finding variants of formulas. Specifically, in this case study, the enumeration helps us to find two safety invariants:
@@ VII-B Specification Repair from Demonstrations
In this use case, we demonstrate how custom optimization objectives can be helpful in learning. The use case is motivated by the field of autonomous agent interpretability using LTL specifications[34]. Consider, for example, the setup in Figure 8a where the end effector of a robotic arm has to be actuated to perform some tasks. The designer provides an initial LTL specification: $\mathbf{F}(green)\land\mathbf{G}(\neg red)$; i.e., reaching the green region and avoiding the red region.
However, after a controller has been developed, the designer decides to refine the LTL specification to better reflect the system requirements. For example, consider the requirement that the robot should satisfy the above LTL formula but also visit the blue region in Figure 8a. The designer can generate a set of demonstrative traces and mark each one as positive or negative, depending on whether the robot successfully reaches the blue region in the trace. Moreover, the modification to the existing LTL should ideally be minimal while taking this additional requirement into account. We pose this as an LTL learning problem, where the associated objective is to minimize the edit distance, i.e., minimizing the removal of existing sub-formulas as well as the addition of new sub-formulas. The encoding of this minimal modification problem is defined as follows:
@@
Lines (8)-(10) define the original formula. Line (11) minimizes the edit distance by using the $max$ operator (with priority 2). The learning process will first retain as many sub-formulas as possible from the old specification and then minimize the total size of the final formula. With these constraints, our approach learns the ideal specification: $\mathbf{F}(green)\land\mathbf{G}(\neg red)\land\mathbf{F}(blue)$, where $blue$ stands for reaching the blue region.
@@ VII-C Invariant Weakening
For instance, Figure 8b is a radiation therapy machine similar to Therac-25 [40] as described in [41]. The machine has two modes: Electron Beam mode and X-ray mode. A spreader must be inserted during the X-ray mode to attenuate the effect of the high-power X-ray beam and limit possible overdose. This safety requirement can be captured as the following invariant $P$: $\mathbf{G}(\textsf{XrayMode}\rightarrow\textsf{SpreaderIn})$.
@@
We consider a safety invariant of the form $\mathbf{G}(\phi\rightarrow\psi)$, where $\phi$ and $\psi$ are Boolean formulas. To achieve weakening, one can learn an invariant $\mathbf{G}(\phi\land\phi^{\prime}\rightarrow\psi\lor\psi^{\prime})$ where $\phi\land\phi^{\prime}$ is a Boolean formula in CNF and $\psi\lor\psi^{\prime}$ is a Boolean formula in DNF. In other words, an invariant can be weakened by (1) adding more conditions (as conjunctions) in the assumption, or (2) adding new acceptable conditions (as disjunctions) in the guarantee. This weakening objective can be expressed through the following constraints:
@@
Specifically, line (12) defines the structure $\mathbf{G}(\phi\rightarrow\psi)$. Line (13) stipulates that $\phi$ and $\psi$ contain only ,,$\land$,$\lor$,$\neg$ , ∨ , ¬, and $AP$. Lines (14)-(16) state that $\phi$ is in CNF and $\psi$ is in DNF. Lines (17)-(18) state that $\phi$ may add conjuncts to XrayMode, and $\psi$ may add disjuncts to SpreaderIn. With these constraints, we can learn the ideal weakened formula as described above.
@@ VIII Evaluation
    RQ1: How does our approach perform compared to the state-of-the-art tool, Flie [9], over unconstrained LTL learning problems?
@@
    RQ2: For problems with structural constraints and/or optimization objectives as described in our case studies, how does our approach perform compared to the state-of-the-art tools?
For RQ1, the metric we are concerned with is the total time taken to find the first satisfying formula. For RQ2, we evaluate performance through two key metrics: (1) the number of constrained problems solved and (2) the time taken to solve them. In particular, even though existing tools cannot encode constraints such as ours, they can be configured to enumerate all solutions until a solution satisfying the constraints is found; the goal here is to compare how long it takes for ATLAS and the baselines to find an ideal solution.
@@ VIII-A Experimental Setup
For fair comparison, we compare ATLAS against the state-of-the-art tools satisfying the following two criteria: (1) no restrictions on the type of formula that can be learned and (2) the guarantee of finding minimal formulas.
@@ VIII-A2 RQ2
For constrained problems, we first compare our tool against Flie, which is configured to enumerate solutions until an expected formula is found. We also compare against LTLSketcher [8] that can learn a formula given a user-defined template. It considers three types of placeholders in a template: (1) any LTL sub-formula, (2) any unary operator, and (3) any binary operator. However, it cannot express all our expected constraints and does not support enumeration either. Thus, we test whether it finds an expected formula in one run with the closest template regarding our required constraints. If it does not, we count it as a timeout. Also note that LTLSketcher is not used in RQ1 as it builds on Flie, and without user-provided templates, it reduces to Flie.
@@ VIII-B Results
Moreover, Table I shows the number of solved problems and the average solving time for Flie, LTLSketcher, and ATLAS. Our tool solves the most number of problems and also has the lowest average solving time. On average, we solve about 3.2x as many problems as Flie and 8.7x as many as LTLSketcher. Also, our tool is about 3x faster than both Flie and LTLSketcher.
Although ATLAS also failed to solve some problems due to timeout, it is better than Flie and LTLSketcher in that: Flie needs to enumerate a large, unconstrained search space which leads to timeout more often. LTLSketcher failed due to the lack of expressiveness (in defining the expected constraints) and enumeration capability. Thus, we show that strong expressive power and the enumeration capability are both critical to LTL learning, and our tool supports both. We can solve all problems that Flie and LTLSketcher can without the loss of generality and much performance overhead.
@@ IX Related Work
Before Neider’s approach, techniques for mining LTL, such as [13, 4, 12], rely on patterns or templates. For example, given a template $\mathbf{G}(x\rightarrow\mathbf{F}y)$, Texada [4] infers specifications by substituting $x$ and $y$ with atomic events from the given traces (and does not learn from negative traces). Compared to our approach, Texada lies on a different trade-off point between expressiveness and efficiency: Although ATLAS gives the user more control over the properties of LTL expressions to be learned, Texada is able to mine from much larger traces (e.g., thousands of events). More recently, Lutz et al. present LTLSketcher [8], where they consider three types of substitutions (i.e., any formula, unary operator, and binary operator) and propose a SAT-based method to find valid substitutions. Although they have improved expressiveness compared to Texada, they cannot express many patterns (e.g., CNF and DNF) as demonstrated in our use cases. In addition, none of these tools supports custom optimization objectives.
Researchers have also explored different variations of the learning problem [47]. Gaglione et al. [48] present an approach using MaxSAT to learn from noisy data. Although $\textnormal{Alloy}^{\textnormal{Max}}$ also uses MaxSAT, we cannot solve this type of problem because $\textnormal{Alloy}^{\textnormal{Max}}$ cannot directly control the weights of clauses in a MaxSAT instance. Roy et al. [49] present an approach for learning from positive-only examples. While our tool also supports such a problem, we can only constrain and optimize the syntactic structure of a solution whereas they can guarantee semantic minimality (i.e., a minimal set of behavior).
@@ X Conclusion and Future Work
We have proposed the constrained LTL learning problem as a generalization of LTL learning from examples, and demonstrated the applicability and performance of ATLAS, our prototype implementation tool. While our evaluation shows promise, we plan to investigate further ways to improve the learning method. First, the scalability of our tool is limited by the underlying MaxSAT solver, and an interesting future work is to investigate an approach that uses a different learning method (based on machine learning, for example). We also plan to investigate methods or heuristics for further reducing the space of possible LTL solutions (e.g., an encoding scheme that eliminates redundant, equivalent LTL expressions). Moreover, we plan to investigate additional use cases for ATLAS (such as invariant synthesis for distributed systems).
Our approach improves the expressive power of LTL learning tools. However, crafting constraints in FOL can be challenging for non-experts in formal logic. As future work, we plan to improve the usability of LTL learning by building on ATLAS (e.g., through a DSL front-end or an LLM that converts natural language to FOL). In addition, ATLAS and other learning tools require the user to provide examples. Future work could explore integrating external methods (e.g., model checkers or test generation tools) to (semi-)automate the generation of examples.
@@ Acknowledgment
This work was supported in part by the NSF awards 2144860 and 2319317, and the NSA grant H98230-23-C-0274. Any views, opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the organizations.